how do you find a equation for a line passing through (3,7) when y varies directly with x

Answer by stanbon(29603)

how do you find a equation for a line passing through (3,7) when y varies directly with x
-------------------------------
y = kx
Find "k":
7 = k*3
k = (7/3)
-----------------------------------
Equation:
y = (7/3)x
==========================
Cheers,
Stan H.